I am so confused on (ln) & (logs) & (e)....
problem
10^(t^2-t)=100
what does this mean? and how do i solve?

Answer by RAY100(1637)

10^(t^2-t) = 100 = 10^2
.
This can be solved without logs by setting exponents equal
.
t^2-t=2
.
t^2 -t -2 =0
.
(t-2)(t+1) =0
.
t= 2, -1
.
check
(t=2),,10^2 = 100,,,,ok
(t=-1),,,,10^2 = 100,,,,ok